OpenStudy (anonymous):

Precal help. A pair of parametric equations is given. x = 2 cos t, y = 3 sin t, 0 ≤ t ≤ 2π

OpenStudy (anonymous):

(a) Sketch the curve represented by the parametric equations. It is an elipse (b) Find a rectangular-coordinate equation for the curve by eliminating the parameter. The answer is \[x^2+\frac{ 4 }{9 }y^2=4\] How do you get this?

ganeshie8 (ganeshie8):

x = 2 cos t => x/2 = cost

ganeshie8 (ganeshie8):

y = 3 sin t => y/3 = sint

ganeshie8 (ganeshie8):

next pull the well known trig identity :sin^2t + cos^2t = 1

ganeshie8 (ganeshie8):

(x/2)^2 + (y/3)^2 = 1

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

Thats not exactly how I got mine x=2cost, y=3sint square both sides x^2=4cos^2t y^2=9sin^2t y^2=9(1-cos^2t) y^2=9-9cos^2t ((y^2-9)/-9)=cos^2t I get x^2=4((9+y^2)/9) Now what I don't get is how they get this to \[x^2= 4-\frac{ 4 }{ 9 }y^2\] which turns into \[x^2+\frac{ 4 }{ 9 }y^2=4\]

ganeshie8 (ganeshie8):

there was a mistake in ur calculation

ganeshie8 (ganeshie8):

Thats not exactly how I got mine x=2cost, y=3sint square both sides x^2=4cos^2t y^2=9sin^2t y^2=9(1-cos^2t) y^2=9-9cos^2t ((y^2-9)/-9)=cos^2t I get x^2=4((9+y^2)/9) ********************************** Now what I don't get is how they get this to

ganeshie8 (ganeshie8):

that line, how come (y^2-9)/-9 turned into (9+y^2)/9 ??

OpenStudy (anonymous):

Negative divided by negative = positive right? This is a method used by my teacher, so I never completely understood how he got that.

ganeshie8 (ganeshie8):

lets see :) you have : \(\large \frac{y^2-9}{-9}\)

ganeshie8 (ganeshie8):

multiply top and bottom wid "-1"

ganeshie8 (ganeshie8):

\(\large \frac{(y^2-9)\times (-1)}{-9\times (-1)} \)

ganeshie8 (ganeshie8):

\(\large \frac{-y^2+9}{9} \)

ganeshie8 (ganeshie8):

\(\large \frac{9-y^2}{9} \)

OpenStudy (anonymous):

Oh Sorry that was what he put, the 9-y^2/9

OpenStudy (anonymous):

So now how does that turn into \[4-\frac{ 4 }{ 9 }y^2\]

ganeshie8 (ganeshie8):

so, you're fine till below line : \(\large x^2=4\left(\frac{9-y^2}{9}\right)\) ?

OpenStudy (anonymous):

Yeah, I don't get how it turns to what I previously showed.

ganeshie8 (ganeshie8):

first divide the 9 inside parenthesis

ganeshie8 (ganeshie8):

\(\large x^2=4\left(\frac{9-y^2}{9}\right)\) \(\large x^2=4\left(\frac{9}{9} - \frac{y^2}{9}\right)\)

ganeshie8 (ganeshie8):

still fine ?

OpenStudy (anonymous):

So that would get you \[x^2=4(1-\frac{ y^2 }{ 9})\] =4-4y^2/9?

OpenStudy (anonymous):

AH alright, that makes total sense. Thank you so much man.

ganeshie8 (ganeshie8):

np.. u wlc :)